In order to estimate the population proportion with a high school education, you need a minimum sample size to result in a margin of error of 7%. For a 95% CI, what is that sample size?
Answer #1

Let proportion, p = 0.50

Margin of error, E = 0.07

Confidence Level, CL = 0.95

Significance level, α = 1 - CL = 0.05

Critical value, z = NORM.S.INV(0.05/2) = 1.9600

Sample size, n = (z² * p * (1-p)) / E² = (1.96² * 0.5 * 0.5)/ 0.07²

= 195.9928 = 196